3sin^2x-4cos^2x belongs to They have given this plz can u give answer to me

Grade:10

1 Answers

Arun
6 years ago
Dear Vishwas
 
3 sin^2 x – 4 cos^2 x = 3 – 7 cos^2 x
 
minimum value = 3 – 7 = – 4
maximum value = 3 – 0 = 3
 
Regards
Arun (askIITians forum expert)